Ask your own question, for FREE!
Biology 12 Online
OpenStudy (anonymous):

Two completely normal parents have a child with a recessive autosomal disorder. How does this happen, and what are the chances?

OpenStudy (alphadxg):

well for this to happen, both parents much represent at least one of the recessive allels, because if one parent did and the other didn't then there is no way that the child would have this disorder. Let me show you what I mean. |dw:1415824902436:dw| The only way for a recessive disorder to be present in the child is the parents must both have a recessive allele for that gene and they most both contribute the recessive allele to b "pp" so the chances are 1/4 = 25% of the child getting this disorder. Please reward a medal if this helped :)

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
1 day ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
1 day ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!